The gas equation is PV = nRT. If P the pressure is 2.0 atm, V the volume of the gas is 6 m3, R the gas constant is 0.08206 m3 atm mole-1 K-1 and T is 300 degrees Kelvin, what are the units of n and what is its numerical value? 


1
Expert's answer
2021-04-14T07:43:12-0400

Q180916

The gas equation is PV = nRT. If P the pressure is 2.0 atm, V the volume of the gas is 6 m3, R the gas constant is 0.08206 m3 atm mole-1 K-1 and T is 300 degrees Kelvin, what are the units of n and what is its numerical value? 


Solution:


please check the value of gas constant, R in the question again


I will first solve the problem using the information given in the problem.

Next I will solve the problem using correct value of 'R' .


Pressure, P = 2.0 atm.


Volume, V = 6.0 m3.


Gas constant, R = 0.08206 m3.atm/mol.K


Temperature, T = 300 K


plug all this information in the ideal gas equation, PV = nRT.


"PV = nRT"

"2.0 \ atm * \ 6.0 \ m^3 = n \ * \ 0.08206 \ m^3.atm/mol.K \ * \ 300K"

"12.0 \ m^3 \ atm = n \ * \ 24.618 m^3.atm/mol"


"n = \frac{12.0 \ m^3 \ atm}{24.618 m^3.atm/mol } = 0.487 \ mol."


The units of n is 'mol' and the numerical value is 0.487.


---------------------------x----------------------------------------------------x-----------------------------------------------------------

I will suggest you to check the question again.

The proper value of Gas constant ' R' is 8.206 *10-5 m3.atm/mol.K


Now I will solve the problem, using R = 8.206 *10-5 m3.atm/mol.K


Pressure, P = 2.0 atm.


Volume, V = 6.0 m3.


Gas constant, R = 8.206 *10-5 m3.atm/mol.K


Temperature, T = 300 K


plug all this information in the ideal gas equation, PV = nRT.


"PV = nRT"

"2.0 \ atm * \ 6.0 \ m^3 = n \ * \ 8.206 *10^{-5} \ m^3.atm/mol.K \ * \ 300K"

"12.0 \ m^3 \ atm = n \ * \ 0.024618 m^3.atm/mol"


"n = \frac{12.0 \ m^3 \ atm}{0.024618 m^3.atm/mol } = 487 \ mol."



If you use the correct value of Gas constant 'R' then we will get

The units of n is 'mol' and the numerical value is 487.